OpenStudy (anonymous):

The legs of a right triangle are 4cm and 8cm long. Find the length of the altitude to the hypotenuse.

OpenStudy (anonymous):

hypotenuse \[\sqrt{4^{2}+8^{2}}=4\sqrt{5}\] \[4*8/2=4\sqrt{5}*a/2\] a=\[8\sqrt{5}/5\] because area of right triangle equals two leg multiplied or hypotenuse multiplied by its altitude